OpenStudy (anonymous):

simplify by rationalizing the denominator: sqrt 2-3/5-sqrt 2

OpenStudy (isaiah.feynman):

\[\frac{ \sqrt{2}-3 }{ 5 - \sqrt{2} }\]. Is that it?

OpenStudy (anonymous):

yes @Isaiah.Feynman

OpenStudy (isaiah.feynman):

So you multiply the fraction by the conjugate of the denominator. Can you do that?

OpenStudy (anonymous):

no @Isaiah.Feynman

OpenStudy (isaiah.feynman):

 Can you go from there?